OpenStudy (anonymous):

When there are 30 girls in a dance, there are 12 boys in the dance. When there are 45 girls in the dance, there are 17 boys in the dance. Assume the relationship is linear. Which equation models the relationship between the number of girls, x, and the number of boys, y? HOLD ON AND I'LL PUT THE ANSWER CHOICES!

OpenStudy (anonymous):

y = 3x + 26 y=1/2x+22 y=1/3x+2 y=3x+9

OpenStudy (anonymous):

@mathmath333 @mathstudent55

OpenStudy (anonymous):

@abb0t @robtobey @Destinymasha

OpenStudy (anonymous):

with every 15 girls there is 2 boys so

OpenStudy (anonymous):

@dumbo So which answer choice would that be?